(a) How many different positive three-digit whole numbers can be formed from the four digits 2, 6, 7, and 9 if any digit can be repeated?

(b) How many different positive whole numbers less than 1000 can be formed from 2, 6, 7, 9 if any digit can be repeated? 

(c) How many numbers in part (b) are less than 680 (i.e. up to 679)? 

(d) What is the probability that a positive whole number less than 1000, chosen at random from 2, 6, 7, 9 and allowing any digit to be repeated, will be less than 680?


1
Expert's answer
2021-05-28T09:52:24-0400

a) Repetitions are allowed

"n=4, r=3"


"n^r=4^3=64"

b) Repetitions are allowed

"n=4, r=1"


"n^r=4^1=4"

"n=4, r=2"


"n^r=4^2=16"

"n=4, r=3"


"n^r=4^3=64"

Then


"4+16+64=84"



c) Numbers from "692" to "699"


"692, 696, 697, 699"

Numbers from "722" to "799"

"n=4, r=2"


"n^r=4^2=16"


Numbers from "922" to "999"

"n=4, r=2"


"n^r=4^2=16"

Then numbers less than 680


"84-(4+16+16)=58"

d)

"P(<680)=\dfrac{58}{84}=\dfrac{29}{42}\approx0.690476"